OpenStudy (anonymous):

3+3x = 5+2x what is x?

OpenStudy (jack1):

3+3x = 5+2x first, minus 3 from both sides of the equals sign write ur answer below dude

OpenStudy (anonymous):

2

OpenStudy (rina.r):

3-5=2x-3x -2 -1x x=-2/-1 x=2.

OpenStudy (jack1):

ur right @vinurules , but better to get the OP to do the steps so they understand it...

OpenStudy (anonymous):

3+3x=5+2x 3x-2x=5-3 x=2 therefore the answer is 2 :)

OpenStudy (shamil98):

Quick find the derivative. \[\large \frac{ x^2 + e^x }{ x^2 - e^x }\]

OpenStudy (anonymous):

u right shamil98 ..but it will be better if u answer it in a easily way :)

OpenStudy (shamil98):

I mean integrate this. \[\large \int\limits_{}^{} xe^{6x} dx\]